A Complete Bibliography of Publications in Spontaneous Generations: A Journal for the History and

Philosophy of Science

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

09 April 2022 Version 1.01

Title word cross-reference

1845 [Heh10]. 18th [Sch18]. 1920 [Gro09]. 1950s [Fra13]. 1950s-1970s [Fra13]. 1970s [Fra13, Auf22]. 19th [Gol16].

2 [Sch18, Thi13b]. 2.0 [Tyf13]. 2005 [RM08]. 2013 [Lei16]. 2014 [Fre16]. 2015 [Fra18]. 20th [Ase08, Ham09].

354pp [Fra18].

6th [SMCL12].

= [Ful07].

A. [Lei16]. Aaron [Gav10]. Abi [Bed16]. Abi-Rached [Bed16]. Account [Kul10]. Acoustical [Pat10]. Action [Sch18]. Activity [Tur13]. Actor [RPK11]. Actor-Network-Theory [RPK11]. Adaptive [Low12]. Adrian [Moo10]. Advertising [Tuc12]. Again [Lad18]. Age [Aga10, Del09, Tur13]. Agenda [Bir13]. Ages [Mil11, Oak18a]. Aggregation [Tur07]. Air [Ott09]. Alexandra [Fra13]. Algorithmic [Poo22]. Allegedly [Heh10]. Alley [RM08]. Allhoff [Vir08b]. Allow [Ros07]. Altered [Pat10]. Alternative [Lyo18]. Ambiguous [Dur12]. America [Gol16]. American [AR11, Ear11, Gro09, Nie09, Tuc12].

Among [Cha16]. Analysis [Byc09]. Anatomical [Ste12, dC07]. Ancestors [Ass10]. Andre [Sch18]. Ann [Wal07]. Anthropological [Bau09]. Anti-realism [Wra18]. Antiquity [Obr12]. Antirealism [Rou18]. Apes [Ass10]. Apparatus [Lev10]. Approaching [Ase08, Bau09]. Approximate [Sta18]. Archaeology [Dis12]. Architectures [Low12]. argument [Psi18]. Aronowitz [AR11]. Art [Tuc12]. Artifacts [Vir10]. artistic [Myb22]. Aspects [Thi13a]. Assault [Gel22]. Asset [HHH22]. Astrolabe [Heh10]. Astronomy [Low12]. Atoms [Mil08]. Authenticating [Heh10]. Avatars [Mil08].

B [Fra13]. Babies [vT07]. Bainbridge [Pet10]. Baldwin [Oak18b]. Bas [For09]. Basalt [Bar10]. Basic [HHH22]. be [Wri09]. Beak [Kem12]. bearers [Per12]. Begin [Jam08]. Behavior [Mil11]. Behaviour [Fra13]. Being [McK18]. Best [Sch22]. Between [Gel22, Lan11, Lev09, Obr12]. BEWARE [Mus18]. Beyond [Rou18, Fra13]. Bias [Poo22]. Big [Men16]. Biobank [TDFM13]. Biodepletion [Aga10]. Biography [Can11]. Biological [Gro09]. Biology [Sch08c]. Blueprints [Ran11]. Blute [Hac09, Wal08]. Bodies [Rob16]. Body [dC07]. Botany [Sch08a]. Boundaries [GRH09, Goo09, Leh09, Lev09, Ott09]. Bowker [Sch07]. Box [Fra13]. Brain [Bed16, Mor09, Sha18]. Brainbound [The09]. Breast [AR11]. Bridge [Wes09]. Bridging [Vir08a]. Brown [Stu12]. Bruce [Aga10]. Bruno [RPK11]. Buchler [Nie09]. Build [Nay22]. Building [Hul07]. Burning [MH12]. Business [Men16].

C [Sch07, Gol16]. C. [Hul07, MN07]. Cambridge [Sch08b]. Can [Jam08, Lor22]. Canada [Auf22, Vir10]. Canadian [Hul10a, Hul10b]. Cancer [AR11, Can11].

Cannot [Ros07]. Can’t [Poo22]. Care [Aga07, Gav10]. Career [Nay22]. Carl [Mor09]. Case [Kem12, MW07, Mil11, Sha18]. Cass [Fre16]. Catalog [Fra18]. Causes [Gol16]. Cautionary [Lyn09]. CE [Heh10]. Cell [Rob16]. Cells [Mai11]. Century [Ase08, Ham09, Pat10, Bar12, Men16, Obr12, Sch18, Ste12]. Ceruzzi [RM08]. Challenge [Heh10, McK18]. Challenges [Lyo18, Ara18]. Challenging [Auf22]. Change [Aga10, McK18]. Changes [Hum10]. Characteristic [HHH22]. Charles [Ear11]. Chemistry [Hen18, The09]. China [Ase08, Chi07]. Chinese [McC10]. Choices [Auf22]. Christianity [Ass10, Jam08]. Christine [Sch08b]. Circle [Ran11]. Civilization [Pet10]. Climate [Aga10, Kel11]. Cluster [Vir10]. Co [Lan11]. Co-operation [Lan11]. Code [Hul07]. Cognition [The09]. Cognitive [Fee10]. Cohen [Gav10]. Cohen-Gadol [Gav10]. Cold [Thi16]. Collection [WD10]. Collections [Heh10]. Colonial [vT07]. Colorblind [Owe16]. Colour [MH12]. Combinatorial [Sch08a]. Comment [AHT09]. Commentary [Wal08]. Commercialization [TDFM13]. Commodification [Ful13]. Complaint [Ear11]. Computational [Spe12]. Computer [Tur13]. Computers [Par10, Poo22]. Concepts [Fee10]. Conceptual [Hum10]. Conjuring [Mye08]. Connecting [Cha22]. Conservative [Jam08]. Conspiracy [Fre16]. Constitution [Low12]. Constraints [Aga07]. Contemporary [Wra18]. Context [Nug16]. Contexts [Var11]. Contingencies [TS09]. Control [Auf22]. Controversial [Boc07]. Controversy [For11, Lig11, Mai11]. Convergences [TS09]. Conversation [Jam08]. Coordination [Pet11]. Copy [Kus12]. Corner [RM08]. Corpse [Ste12]. Cosmology [Pri07]. Cover [Ano07, Ano08].

Craft [Lan11]. Craver [Mor09]. Create [Mai11, Men16]. Creation [Leh09, Wes09]. Crisis [Eya22, Aga10]. Crist [Aga10]. Critical [Thi13b]. Criticisms [Ala18]. Crossing [TS09]. Cubism [Myb22]. Cuckoo [Pri07]. Cultural [Bau09]. Culture [Bau09, Rec10, SJ10, Sis22, Vir10]. Cultures [Lev09]. Cures [Joh22b]. Curie [Goo09]. Current [Vir08b]. Cushing [Gav10]. Cyberscience [Sch08b]. Cynicism [Lyn07].

D [Gav10]. Dangerous [Fre16]. Daniel [Rec09]. Dark [Mil11]. Database [Hen12, Fra18]. Dated [Heh10]. David [Pat10, Thi13b]. Deafness [Joh22b]. Dealing [Wal09]. Debate [Sha18, Wra18, Ass10]. Decommissioned [Ful07]. Deep [Lor22]. Defence [Ful13]. Definition [Blu08, Wal08]. Deliberations [Var11]. Democratic [Gre22]. Democratised [Ful07]. Dennis [Gav10]. Deployment [Ala18]. Depth [Kul10]. Descriptive [Sch08a]. Design [AHT09]. Development [Nug16, Shi10]. Devo [Blu08]. Diagram [MH12]. Diagrams [Dur12, Wüt12]. Digital [Par10]. Dilemma [San18]. Dilemmas [Mil07a]. Disagreement [Wal09]. Disappear [Fla12]. Disappearing [Bur11]. Disaster [Var11]. Disciplinary [Sen13]. Disciplining [Vis08]. Discussion [Mil07a]. Disease [Chi07]. Dissection [Ste12]. Distinguishing [Wal09]. Divergences [Eps16]. Diversity [Owe16]. Do [Dow12, Kel11, Mai11]. Dobsonian [SJ10]. DOG [Mus18]. Domesticating [Goo09, Shi10]. Door [WD10]. Doppelt [Ala18]. Dorothy [Men16]. Douglas [Lei16, MN07, Oak18a]. Dowd [Oak18a]. Drake [Oak18a]. Dreams [Fra18]. Drugs [Sis22]. Dunlap [Cha16]. Dynamics [Goo12].

Early [Goo09, Ham09, Ase08, Obr12, Ste12].

Earth [Aga10, Lor22]. Earthquakes [Var11]. Eco [Blu08]. Economic [MN07, Thi13a]. Economics [Blu13, Sen13, Thi13b, TDFM13]. Economy [Bir13, Tyf13]. Editing [Ros22]. Editor [GL12, For11, For18, GRH09, Mil07a, Rec08, Rec10, Thi13a, WW16]. Eds [Vir08b, Aga10, Sch18]. Education [AHT09, Ste12]. Effective [Boc07]. Eileen [Aga10]. Einstein [Lei16]. Electrodynamics [Nug16]. Elements [Bur11]. Embers [MH12]. Emergence [Gro09]. Emergent [Wes09]. Emerging [Bir13, Sch08c, Vir08b]. Emperor [Can11]. Engagement [Ros22, Myb22]. Engaging [Ara18, TN10]. Engine [Hul10a, Hul10b]. Engineer [MN07]. Engineering [Bai10]. Environment [Boc07]. Environmental [Cha22]. Epistemic [Alm07, GRH09, Goo09, Leh09, Lev09, Lyo18, Nie09, Ott09, Wal09]. Epistemological [Mil07a]. Epistemology [Tur07]. Equation [Oak18a]. Equipment [Har10]. Estimating [Oak18a]. Ethical [Vir08b]. Ethics [Aga10, Aga07, Sch08c]. Euclidean [Bar12]. Eugenics [Dyc22]. European [SMCL12]. Every [Del09, Wri09]. Everyday [Lew12]. Evidence [Alm07, Bue08, Wra18]. Evo [Blu08]. Evolution [AHT09, Ass10, Blu08, Jam08, Wes09, Wes09]. Evolutionary [Blu13]. Examining [Vir08a]. Exceeding [Del09]. Exchange [Bau09]. Exhausting [Ase08]. Expectations [MW07]. Experiment [Har10]. Experimental [Byc09, Fee10]. Experimentation [Lev10]. Experiments [Wal07]. Expert [Hul07, Low12, Nay22, Ott09]. Expert-Resident [Ott09]. Expertise [Cha22, Dyc22, Eya22, Ful07, Gel22, HHH22, Lyn07, Mil07a, Reh22, Ros22, dC07, vW22]. Experts [Alm07, Gre22, Tur07]. Explaining [Mor09]. Explanation [Kul10]. Exploration [MH12]. Exploring [Lan11, Lev09]. Exposing [Heh10].

Extended [Cha10, The09]. Extraterrestrial [Oak18a]. eyes [Kem12].

F [Fra13, Mor09, Oak18a]. Fake [Poo22]. Falsely [Fra10]. Farewell [Sta18]. Farmed [Boc07]. Fatal [Men16]. Feathered [Cha16]. Fencelines [Ott09]. Feyerabend [Roz13]. Feyerabendian [Fos18]. Feynman [Wüt12]. Field [Bau09, Cha16]. Fields [Tur13]. Fifth [Rie11]. Figure [Sch18]. Figures [Obr12]. First [Men16]. Flatter [Fla12]. Focused [Mil07a]. Fond [Sta18]. Forgery [Heh10]. Four [Lyo18]. Fraassen [For09]. Framework [HHH22]. Frankenstein [Rec08]. Frederick [Lew12]. Freeness [Hul10a, Hul10b]. Fritz [Vir08b]. Front [Ano22]. Frontier [Rob16]. Fuller [Thi11]. Functional [Sch08a]. Funding [Bro13]. Future [HH18, Lor22].

Gadol [Gav10]. Gaia [Aga10]. Gap [Ros07, TS09, Vir08a]. Gender [Jac08, Wal07]. Gendered [Ham09]. Generating [TN10]. Generation [Fee10]. Genesis [Nug16]. Genetic [Gro09]. Genome [Ros22]. Geoffrey [Sch07]. Geology [Shi10]. Global [Wri09]. Goes [Ful08]. Goethe [Hal09]. Going [Byc09, Lyn09]. Government [MW07]. Gravitational [Col18]. Great [Bar10]. Grinnell [Lew12].

H [Aga10, MN07]. Hands [Hen12, Ols10]. Handservant [Ros22]. Happiness [Joh22b]. Hartlib [Ran11]. Harvey [Gav10]. Haven [Fra18]. Health [Chi07]. Hearing [Joh22b]. Hegemonic [Hul10a, Hul10b]. Helmholtz [Hal09]. Help [Lor22]. Herbarium [Fla12]. Heritable [Ros22]. Hermann [Hal09]. Hermaphroditic [dC07]. Hesketh [Ass10]. Heterogeneous [Bau09]. hidden [Tuc12]. High [RM08]. Hijacking [Moo10]. Hine [Sch08b]. Historical [Low12, Sen13, WD10]. Historiography [The09]. History [AR11, Bur11, Cou10, Dyc22, Hen18, Oak18b, Rie11, SMCL12, WD10, Wra18, Ara18, Psi18, Joh22b]. hold [Wri09]. Holdings [Tay12]. Holenstein [Sch18]. HPS [Rie11]. Hubert [Sch18]. Human [Mil11, Ros22, Sha18]. Humanity [Fra18]. Humans [Col10]. Hygienic [Chi07].

Ian [Ass10]. Ideal [MN07]. Ideas [Gol16, Fre16]. Illustration [Ste12]. Illustrators [Hen12]. Images [Dis12, Dow12, Spe12]. Importance [Nie09, Owe16]. Independence [Alm07]. Independent [Sch08a]. India [Var11]. Inequalities [Eps16]. Inequality [HHH22, Pur16, WW16]. Infant [vT07]. Infinite [Pri07]. Information [Gro09, Poo22, Tur13]. Informationalising [Kea08]. Institute [Vir10]. Instructive [Ste12]. Instrument [Par10]. Instruments [Bun10, Col10, Der10, Ols10, Ott09, Rec09, Rec10, Shi10, Tur13, WD10]. Integrated [Rie11]. Intelligent [AHT09]. Interdisciplinary [Ols10]. International [SMCL12]. Internet [RM08]. Interpreting [Wüt12]. Intertheoretic [Nug16]. Intrinsically [Dur12]. Introduction [For11, For18, GRH09, GL12, Mil07a, Rec08, Rec10, Thi13a, WW16, vT07, RPK11, vW22]. Intuition [Lew12]. Invention [Men16]. Inventions [Sch08a]. Invisible [Hen12]. IPCC [MH12]. Issues [Vir08b].

Jacob [Wes09]. James [Stu12]. Joelle [Bed16]. John [Gol16, Pri07]. Joint [Rie11]. Josef [Reh22]. Journal [Oak18b]. Jurisprudence [Ase08]. Justification [Nie09]. Justus [Nie09].

Kantian [Gel22]. Kean [Bur11]. Kearsley [Gol16]. Kekulé [The09]. Kerala [Var11]. Kirke [Kus12]. Klapwijk [Wes09]. Knowing [Wal07]. Knowledge [Blu09, Cha10, Fee10, Ful13, Rec10, Sch18, Tur07, Vir08a]. Koenig [Pat10].

Laboratories [Mil08]. Laboratory [Fra13, Stu12, Wri09]. Labs [Lan11]. Lange [Bel10]. Last [Rou18]. Laszlo [Myb22]. Late [Cou10, Gol16]. Latour [RPK11]. Law [Ase08]. Lawmakers [Bel10]. Laws [Bel10, Bel10]. Learning [HH18, Ols10, Vir10]. Lee [Mil11]. Legacy [Gav10]. Lemov [Fra18]. Leslie [Pri07]. Lessons [Lyn07]. Let [Hul10a, Hul10b]. Life [Cou10, Fra13, Mye08, Oak18a]. Lilliput [Rec08]. Limb [Gol16]. Limits [Vir08a]. Lin [Vir08b]. Living [Wes09]. Logic [Lew12]. Lost [Fra18]. Love [Bur11]. Lynch [Ful09].

MA [Sch08b]. Machine [Fra10, HH18, TN10]. Machinic [Mye08]. Mad [Jac08, Mus18]. Madness [Bur11]. Magnet [Goo09]. Make [Fla12]. makers [Per12]. Making [Car12, Hen12, Wri09, Oak18b]. Maladies [Can11]. Mammal [Kem12]. Management [Bed16]. Managerialism [Gre22]. Managing [MW07]. Manifold [Poo22]. Maó [SMCL12]. Marc [Bel10]. Maria [Ham09]. Marie [Goo09]. Martin [Sch18]. Massively [Mil08]. Massively-Multiplayer [Mil08]. Material [Bar10, SJ10, Vir10]. Materials [Ham09]. Maternal [vT07]. Matter [Ano22, Beu12, Kea08]. Matters [Sha18]. Matthew [Oak18a]. Maxwell [TS09]. Maxwellian [Nug16]. May [SMCL12]. McIntyre [Mil11]. Meanings [Chi07]. Measuring [Bun10]. Mechanisms [Mor09]. Media [Var11]. Mediators [Ols10]. Medical [Aga07, Ste12]. Medicine [Ear11]. Meet [Lew12]. Melinda [Oak18b]. Memory [Sch07]. Menorca [SMCL12]. Metaphor [Pit08]. Metaphysics [Bel10, McK18]. Method [Wal07]. Metrology [Bar10]. Michael [Kul10]. Microfluidics [Ols10]. Microscope [Tur13].

Mike [Ful11]. mimic [Car18]. Mind

[The09, Bed16, Stu12]. Minds

[Pri07, Rec09]. Mistrust [Eya22]. Mitchell [Gol16]. Model [Byc09]. Modeling [Goo12]. Models [Win08, Wüt12]. Modern [Cou10]. Modernity [Chi07]. Moholy [Myb22]. Moholy-Nagy [Myb22]. Monitoring [Ott09]. Moral [Cou10]. Moroccan [Heh10]. Mosaic [Mor09]. Most [HHH22]. Mothers [vT07]. Motives [Fin18]. Much [Dow12, Gol16]. Mukherjee [Can11]. Multiplayer [Mil08]. Multiple [Sch08a]. Museum [Heh10, Vir10]. my [Kem12]. Myths [Leh09].

Nagel [AHT09]. Nagy [Myb22]. Names [Sch08a]. Nano [Vis08, Win08]. Nano-scale [Win08]. Nanomedicine [Vir08a]. Nanoscale [Bue08, Kea08, Rec08]. Nanotechnologies [Beu12]. Nanotechnology [Jac08, Pit08, Sch08c, Vir08b]. Natural [Blu08]. Naturalism [Nie09]. Nature [Dur12, Tuc12, Car18, Oak18b, Bel10]. Naval [Wri09]. navigator [Wri09]. Network [RPK11]. Neuro [Bed16]. Neuropsychology [Fee10]. Neuroscience [Mor09]. News [Poo22]. Newton [TS09]. Nigeria [vT07]. Nikolas [Bed16]. Nineteenth [Pat10, Ste12]. Nineteenth-Century [Pat10, Ste12]. Non [Das09, Sch08a]. Non-Functional [Sch08a]. Non-Western [Das09]. Notes [Myb22]. Nowhere [Dis12]. Numbers [Eya22].

Oakley [Wal07]. Objectivity

[Hal09, Lew12]. Objects

[Del09, Heh10, Nie09]. Observatory [Wri09]. Old [Ols10]. Ongoing [Ful08]. operation [Lan11]. Optics [Low12]. Ordinal [Nie09]. Other [Bur11, Fre16]. Otto [McC10]. our [HHH22, Ear11]. Overview [Thi13b]. own [Car18, Kem12]. Oxford [Ass10].

Page [Ano07, Ano08]. Pain [Gol16]. Pancake [Fla12]. Pantalony [Pat10]. Paradoxes [For09]. Paris [Pat10]. Parr [Moo10]. Passion [Lew12]. Past [Dis12]. Patient [Gav10]. Patrick [Vir08b]. Paul [RM08, Roz13]. Pedagogy [Jam08]. Pedantry [Gel22]. Peer [Ran11]. Peers [Wal09]. People [Der10, Rob16, Cha18]. Perception [Low12]. Perceptions [Owe16]. Performance [Ful08]. Periodic [Bur11]. Periodicals [Lig11]. Peripheries [Das09]. Permanency [Goo09]. Personal [Nay22]. Perspectival [Tel18]. Perspective [Gel22, Sen13, For09]. Phantom [Gol16]. Phenomenology [Nie09]. Phillip [Gro09]. philosopher [Wri09]. Philosophical [Pri07, Reh22, Rec09, Kus12]. philosophically [Ara18]. Philosophizing [Reh22]. Philosophy [Car12, Jam08, Rie11, The09]. Physics [Nay22, Spe12, Myb22]. Pieper [Reh22]. Placing [Del09]. Planetary [Shi10]. Planets [Shi10]. Point [Bar12]. Policy [Cha22, Sha18]. Political [Bir13, Mil07a, Tur07]. Politics [Bar10, Bro13, Men16]. Popular [Lev09, Sch22]. Popularization [Sch08c, SMCL12]. Population [Auf22]. Populism [Gel22]. Port [Chi07]. Potential [Vir08a]. pp [Fre16, Lei16]. Practice [Lew12, Par10, Pet11, Rec10, Sch18]. Practices [Ham09, Sch07, Shi10, Car18]. Pragmatism [Fos18]. Prehistory [Ran11]. Present [Ear11]. Presented [Vir10]. Press [Fra18, Lei16, Var11, Men16, Sch08b]. Prevalence [Oak18a]. Princeton [Lei16]. Principles [Ase08]. Print [Bar12, Tuc12]. Privatization [Bro13]. Problem [Pet11, TDFM13]. Problematic [MW07]. Professor [Hul07, Hac09]. Profiles [Gav10]. Progress [Das09, Ful13]. Project [MW07, Sha18]. properly [Wri09]. Psychiatry [Dyc22]. Public [AHT09, Bro13, For11, Ful08, Lyn09, Mai11, MW07, Ros22]. Publication [TN10]. Purpose [Wes09]. Pyramid [Bar10].

qualified [Wri09]. Quantum [Lei16]. Quest [Fra18, Lei16]. Quo [Vic18].

R [Cha16, Fre16, Hul07, Stu12]. Race [Men16]. Rached [Bed16]. Racial [Owe16]. Radium [Ham09]. Rationality [Gro09]. Re [Men16]. Re-Create [Men16]. Reading [Bun10, Vir10]. Realism [Ala18, CvF18, Cha18, Col18, Hen18, HH18, Lad18, Lyo18, Rou18, Sha18, Tel18, Vic18, Ara18, McK18, Psi18, Wra18]. Realism/Anti [Wra18]. Realism/Anti-realism [Wra18]. Realist [San18, Thi13b, Mus18]. realistic [Cha18, McK18]. Realities [Bai10]. Reassembling [RPK11]. Rebecca [Fra18]. Reckoning [Lor22]. Referential [Tel18]. Reflections [Blu09, Nay22]. Reforms [MN07]. Regional [Var11]. Regulation [Sis22]. Regulatory [Eya22]. Reinventing [Dyc22]. Relevance [Pet11, Wra18]. Religious [Ran11]. Remoralizing [Tyf13]. Rentetzi [Ham09]. Report [Ful08]. Representation [Beu12, For09, GL12, Obr12]. Representations [Bar12, Goo12, SMCL12]. Research [Bau09, Bir13, Fin18, Ham09, Lan11, Owe16]. Resident [Ott09]. Resource [Aga07]. Response [Ful09, Ful11, Hac09]. Responsibility [Reh22]. Resurgences [Eps16]. Review [AR11, Aga10, Ass10, Auf22, Bed16, Bel10, Bur11, Can11, Cha16, Cha22, Chi07, Cou08, Cou10, Ear11, For09, Fra13, Fra18, Fre16, Gav10, Gro09, Ham09, Joh22b, Kul10, Lei16, Lew12, Lor22, Men16, Mil11, Moo10, Mor09, Oak18a, Oak18b, Pat10, Pet10, Poo22, Pri07, Ran11, RM08, Rec09, Rob16, RPK11, Roz13, Sch07, Sch08b, Sch18, Sis22, Ste22, Stu12, SMCL12, Thi11, Thi13b, Thi16, Vir08b, Vir10, vW22, Wes09, Mil07b, Rie11]. Revisited [Lev10]. Rights [Rob16]. Ring [Hul10a, Hul10b]. Rinker [Aga10]. Risk [MH12]. Robert [AR11]. Roberts [Men16]. Robustness [Byc09]. Rogaski [Chi07]. Role [Lan11]. Room [McC10]. Rose [Bed16]. Rosenberg [Ear11]. Rothbart [Rec09]. Rudolph [Pat10]. Ruth [Chi07]. Rutherford [Fra13].

S. [Gol16]. Sam [Bur11]. Savant [Sch18]. Scale [Beu12, Win08]. Scanning [Fla12]. Scholars [Sch18]. School [SMCL12]. Schuster [Fre16]. Science [Ase08, Bel10, Blu13, Boc07, Byc09, Car12, Cha22, Cou08, Das09, Dis12, Eps16, Eya22, For11, Ful07, Ful08, Gro09, GL12, Jam08, Joh22a, Lew12, Lyn07, Men16, Mil07b, Mil11, Obr12, Owe16, Pet10, Rec08, Rob16, Sch22, Sen13, Sha18, Sis22, Ste22, SMCL12, Thi13a, Thi13b, Tyf13, Var11, Vir10, Wra18, Ara18, Psi18, Tuc12, Rie11, Tyf13, Vir10, Roz13, Thi11, Thi16]. Sciences [Bed16, Wal07, Sch07]. Scientific [Bau09, CvF18, Col18, Cou10, Der10, Dow12, Gel22, Goo12, Hal09, Heh10, Hen18, Hen12, Kul10, Lad18, Lan11, Lev09, Lyo18, Mil07a, Oak18b, Par10, Pet11, Rec10, San18, Vic18, WD10, Ara18, McK18, For09]. Scientists [Jac08, Kel11]. Sea [Hum10]. Search [Cou08]. Secrecy [Goo09]. Secularity [Goo09]. Seeing [Dis12]. Seeking [Boc07]. Selection [Blu08]. Selective [Vic18]. Sensations [Pat10]. Services [vT07]. shape [Bar12]. Shapin [Cou10]. Sheets [Fla12]. Short [WD10]. Shouldn’t [Fla12]. Siddhartha [Can11]. Simon [Fre16]. Sims [Pet10]. Simulation [Par10]. Situating [Cou08, Vir10]. Sixteenth [Bar12]. Sixteenth-Century [Bar12]. Skepticism [Lyn07]. Skinner [Fra13]. Sloane [Del09]. Small [Pit08, Bar12]. Social [Ase08, Eps16, Pet10, Pur16, RPK11, Thi16, Wal07, WW16]. Society [Ful13, MN07, Vir08b, AR11, Cha22]. Socratic [Lyo18]. Solve [Poo22]. Sound [JI12]. Space [Dur12, Gro09, Myb22]. Space-Time [Myb22]. Spain [SMCL12]. Speaks [Fra10]. Specimens [Ste12]. Spencer [Gav10]. Spoon [Bur11]. Spring [SMCL12]. Standard [Hul10a, Hul10b]. Starving [Pri07]. State [vT07]. Steinke [Sch18]. Stem [Mai11, Rob16]. Steve [Thi11]. Steven [Cou10]. Stone [Lei16]. Strange [Kem12, Sis22]. Strevens [Kul10]. Structure [Goo12]. STS [WW16]. Stuber [Sch18]. Studies [Ful08, Lyn07]. Study [MW07]. Studying [Eps16]. Stuttgart [Hen12]. Subject [Hal09]. Such [Mai11]. Suffering [Gol16]. Summer [Vir10]. Sunstein [Fre16]. Supported [Cou08]. Sustainability [Moo10]. Swabian [Lei16]. Synthetic [Sch08c]. Systematics [Sch08b]. Systems [Kea08, MW07].

Table [Bur11]. Tale [Lyn09]. Tales [Bur11]. Talk [Pit08]. Taming [TN10]. Technocracy [Ros22]. Technological [MW07]. Technologies [Sch08c]. Technology [Fra13, Lyn07, Ols10, Pur16, RM08, Sch08a, Vir10]. Technoscience [Bir13]. Telescope [SJ10]. Terminated [Nay22]. Tester [Hul10a, Hul10b]. testimony [Kem12]. Them [Fla12]. Theological [Pri07]. Theoretical [Car18]. Theories [Fre16, Win08]. Theory [Gre22, Lan11, McK18, RPK11]. Thicke [Ful11]. Thing [Cha10, Del09]. Think [Gol16]. Thinking [Lor22]. Thirteenth [Obr12]. Thomas [AHT09, Cha16, Kus12]. those [Car18]. Thought [Bau09]. Thurtle [Gro09]. Time [Blu08, Dur12, Gro09, Lor22, Myb22]. Together [Reh22]. Tolstoy [Psi18]. Tools [Fee10, Rec09]. Toronto [WD10, Hul07]. Trading [TN10]. Trafficking [Ham09]. Trans [Eya22]. Trans-science [Eya22]. Transactions [Kus12]. Transdisciplinarity [Cou08]. Transition [Tyf13]. Treaty [Chi07]. Treaty-Port [Chi07]. Trees [Blu09]. Trips [Sis22]. Trouble [Spe12]. True [Bur11]. Trust [Mil07b, Ste22]. Truth [Per12, Sta18]. Truth-bearers [Per12]. Truth-makers [Per12]. Turmoil [Aga10]. Twenty [Men16]. Twenty-First [Men16]. Two [Cou08, Joh22a, Wal09, Ara18]. Tyfield [Thi13b]. Types [Wal09]. Tyranny [Roz13]. Tysons [RM08].

U.S. [Wri09]. Unbounded [Wes09]. Understandings [Kea08]. Unity [Mor09, TN10]. Universe [Obr12]. University [Bro13, Fra18, Lei16, WD10]. Unnatural [AR11]. unsensible [Bar12]. Updated [Blu08, Wal08]. Utopia [Myb22].

Vadis [Vic18]. Vakoch [Oak18a]. Valiant [Lei16]. Vein [Nie09]. Versus [Tur13]. Vienna [Ham09]. View [MN07]. Virtual [Mil08, Pet10]. Virtues [Joh22a]. Visible [Car12, Hen12]. Vision [JI12]. Visual [Bar12, Beu12, Bue08, Car12, Goo12, GL12, Obr12, Tur13, Wüt12, SMCL12]. Vocation [Cou10]. Volumes [Sch18, Thi13b].

War [Thi16]. Warcraft [Pet10]. Was [The09]. Waves [Col18]. Weir [Gol16]. Welfare [vT07]. Western [Das09]. Where [Lew12]. Wikipedia [Ros07]. Wild [Boc07]. William [Pet10]. within [Lan11]. withstands [Ala18]. Work [Dow12, Goo09, Car18, Rec09]. Workshop [Mil07b, Pat10, Rie11, SMCL12]. Workshops [Cou08]. World [Bur11, Wes09, Tuc12, Pet10]. Worlds [Mil08]. Worries [Mil07a].

Yale [Fra18]. York [Men16]. Young [Hul07].

Zones [TN10].

References

Agassi:2007:EMC

[Aga07] Joseph Agassi. On the ethics of medical care under resource constraints. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1 (1):4, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

                                                  spontaneousgenerations.library.
                                                  utoronto.ca/index.php/SpontaneousGenerations/
                                                  article/view/2966.

Agapitos:2010:REC

[Aga10] Julia Agapitos. Review: Eileen Crist and H. Bruce Rinker, eds. Gaia in Turmoil: Climate Change, Biodepletion and Earth Ethics in an Age of Crisis. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):286–288, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                                  spontaneousgenerations.library.
                                                  utoronto.ca/index.php/SpontaneousGenerations/
                                                  article/view/14192.

Aikin:2009:EID

[AHT09] Scott Aikin, Michael Harbour, and Robert Talisse. Evolution, intelligent design and public education: A comment on Thomas Nagel. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):35– 40, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                                  spontaneousgenerations.library.
                                                  utoronto.ca/index.php/SpontaneousGenerations/
                                                  article/view/6103.

Alai:2018:HDR

[Ala18] Mario Alai. How deployment realism withstands Doppelt’s criticisms. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):122– 135, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27046.

Almassi:2007:EEE

[Alm07] Ben Almassi. Experts, evidence, and epistemic independence. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1): 58, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2972.

Anonymous:2007:CP

[Ano07] Anonymous. Cover page. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):??, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5077.

Anonymous:2008:CP

[Ano08] Anonymous. Cover page. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):??, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.

                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/5076.

Anonymous:2022:FM

[Ano22] Anonymous. Front matter. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):??, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38289.

Abi-Rached:2011:RRA

[AR11] Joelle M. Abi-Rached. Review: Robert A. Aronowitz. Unnatural History: Breast Cancer and American Society. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1): 79–82, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14569.

Arabatzis:2018:EPH

[Ara18] Theodore Arabatzis. Engaging philosophically with the history of science: two challenges for scientific realism. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 35–37, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/27095.

Asen:2008:ALE

[Ase08] Daniel Asen. Approaching law and exhausting its (social) principles: Jurisprudence as social science in Early 20th Century China. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1): 213, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/3511.

Assenza:2010:RIH

[Ass10] Sebastian Assenza. Review: Ian Hesketh. Of Apes and Ancestors: Evolution, Christianity, and the Oxford Debate. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):262–265, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/12020.

Auffrey:2022:RCC

[Auf22] Vincent Auffrey. Review of Challenging Choices: Canada’s Population Control in the 1970s. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):132– 134, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38213.

Baird:2010:ER

[Bai10] Davis Baird. Engineering realities. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):94– 110, ???? 2010. CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11332.

Barany:2010:GPM

[Bar10] Michael J. Barany. Great Pyramid metrology and the material politics of basalt. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 45–60, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11897.

Barany:2012:SUS

[Bar12] Michael Jeremy Barany. “That small and unsensible shape”: Visual representations of the Euclidean point in Sixteenth-Century print. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):148– 159, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16137.

Baus:2009:CEH

[Bau09] Daniela Baus. Cultural exchange in a heterogeneous research field: Approaching scientific culture with anthropological thought. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1): 95–104, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6574.

Bedford:2016:RNR

[Bed16] Riiko Bedford. Review: Nikolas Rose and Joelle Abi-Rached, Neuro: The New Brain Sciences and the Management of the Mind. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):93– 96, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/21160.

Belanger:2010:RML

[Bel10] Christopher Belanger. Review:

Marc Lange. Laws and Lawmakers: Science, Metaphysics, and the Laws of Nature. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):266–269, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/12682.

Beumer:2012:MSV

[Beu12] Koen Beumer. A matter of scale: The visual representation of nanotechnologies. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):65–74, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16139.

Birch:2013:PET

[Bir13] Kean Birch. The political economy of technoscience: An emerging research agenda. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):49–61, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19556.

Blute:2008:ITU

[Blu08] Marion Blute. Is it time for an updated ‘eco–evo–devo’ definition of evolution by natural selection? Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1): 1, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/4896.

Blute:2009:RTK

[Blu09] Marion Blute. Reflections on trees of knowledge. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1): 223–225, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/8474.

Blute:2013:EES

[Blu13] Marion Blute. The evolutionary economics of science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):62–68, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.

utoronto.ca/index.php/SpontaneousGenerations/
article/view/19562.

Bocking:2007:WFS

[Boc07] Stephen Bocking. Wild or farmed? Seeking effective science in a controversial environment. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):48, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2971.

Brown:2013:PUF

[Bro13] Mark B. Brown. Public university funding and the privatization of politics. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1): 21–28, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/20002.

Bueno:2008:VEN

[Bue08] Otávio Bueno. Visual evidence at the nanoscale. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2 (1):132, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5075.

Bunge:2010:RMI

[Bun10] Mario Bunge. Reading measuring instruments. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 85–93, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11725.

Bursten:2011:RSK

[Bur11] Julia R. Bursten. Review: Sam Kean. The Disappearing Spoon, and Other True Tales of Madness, Love, and the History of the World from the Periodic Table of the Elements. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1): 100–102, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14955.

Bycroft:2009:GOM

[Byc09] Michael Trevor Bycroft. Going outside the model: Robustness analysis and experimental science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):123– 141, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6118.

Canavan:2011:RSM

[Can11] Barbara C. Canavan. Review: Siddhartha Mukherjee. The Emperor of All Maladies: A Biography of Cancer. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):103– 105, ???? 2011. CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14963.

Carusi:2012:MVV

[Car12] Annamaria Carusi. Making the visual visible in philosophy of science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):106– 114, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16141.

Cartwright:2018:TPW

[Car18] Nancy Cartwright. Theoretical practices that work: those that mimic Nature’s own. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):165–173, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27045.

Charbonneau:2010:ETK

[Cha10] Mathieu Charbonneau. Extended thing knowledge. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):116–128, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11891.

Charenko:2016:RTR

[Cha16] Melissa Charenko. Review: Thomas R. Dunlap, In the Field, Among the Feathered. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):97–99, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/19962.

Chang:2018:RRP

[Cha18] Hasok Chang. Realism for realistic people. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 31–34, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/27002.

Charriere:2022:REE

[Cha22] Elodie Charrière. Review of Environmental Expertise. Connecting Science, Policy, and Society. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):129– 131, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38212.

Chiang:2007:RRR

[Chi07] Howard Hsueh-Hao Chiang. Review: Ruth Rogaski. Hygienic Modernity: Meanings of Health and Disease in Treaty-Port China. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):146, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2981.

Collins:2010:HI

[Col10] Harry Collins. Humans not instruments. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 138–147, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11354.

Collins:2018:GWS

[Col18]  Harry Collins.  Gravitational

waves and scientific realism. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):38–41, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27042.

Cournoyea:2008:STR

[Cou08] Michael Cournoyea. In search of transdisciplinarity: A review of two workshops supported by situating science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2 (1):238, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5072.

Cournoyea:2010:RSS

[Cou10] Michael Cournoyea. Review: Steven Shapin. The Scientific Life: A Moral History of a Late Modern Vocation. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 273–275, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/13860.

Chakravartty:2018:WSR

[CvF18] Anjan Chakravartty and Bas C. van Fraassen. What is scientific realism? Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 12–25, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/26992.

Dasgupta:2009:PSS

[Das09] Deepanwita Dasgupta. Progress in science and science at the non-Western peripheries. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):142–157, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/6575.

daCosta:2007:AEH

[dC07] Palmira Fontes da Costa. Anatomical expertise and the hermaphroditic body. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):78, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/1016.

Delbourgo:2009:EAE

[Del09] James Delbourgo. ‘Exceeding the age in every thing’: Placing Sloane’s objects. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):41–54, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6743.

Derksen:2010:PSI

[Der10] Maarten Derksen. People as scientific instruments. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):21–29, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11842.

Dissard:2012:SPN

[Dis12]  Laurent Dissard.  Seeing the

past from nowhere: Images and science in archaeology. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):24–33, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16144.

Downes:2012:HMW

[Dow12] Stephen M. Downes. How much work do scientific images do? Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):115– 130, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/17154.

During:2012:IAN

[Dur12] Elie During. On the intrinsically ambiguous nature of space–time diagrams. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 160–171, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/18645.

Dyck:2022:REH

[Dyc22] Erika Dyck. Reinventing expertise in the history of psychiatry and eugenics. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1): 107–112, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38206.

Earl:2011:RCR

[Ear11] Martin J. Earl. Review: Charles E. Rosenberg. Our Present Complaint: American Medicine, Then and Now. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):

83–84, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14712.

Epstein:2016:SSS

[Eps16] Steven Epstein. Studying science and social inequalities: Resurgences and divergences. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):3–12, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/21366.

Eyal:2022:MNR

[Eya22] Gil Eyal. Mistrust in numbers regulatory science, transscience and the crisis of expertise. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):36– 46, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38197.

Feest:2010:CTE

[Fee10] Uljana Feest. Concepts as tools in the experimental generation of knowledge in cognitive neuropsychology. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 173–190, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11938.

Fine:2018:MR

[Fin18] Arthur Fine. Motives for research. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):42– 45, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/27048.

Flannery:2012:FTP

[Fla12] Maura C. Flannery. Flatter than a pancake: Why scanning herbarium sheets shouldn’t make them disappear. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 225–232, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/16134.

Forbes:2009:RBV

[For09] Curtis Forbes. Review: Bas van Fraassen. Scientific Representation: Paradoxes of Perspective. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):236– 238, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/6666.

Forbes:2011:SPC

[For11] Curtis Forbes. Science and public controversy: Editor’s introduction. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):

1–4, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/15622.

Forbes:2018:EI

[For18] Curtis Forbes. Editor’s introduction. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 1–11, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/29354.

Foss:2018:FP

[Fos18] Jeff Foss. Feyerabendian pragmatism. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):26– 30, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/29355.

Franklin:2010:MSF

[Fra10]  Allan Franklin.  The ma-

chine speaks falsely. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):71–84, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11333.

Fraser:2013:RAR

[Fra13] Jennifer Fraser. Review: Alexandra Rutherford, Beyond the Box:

B. F. Skinner’s Technology of Behaviour from Laboratory to Life, 1950s-1970s. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1): 100–102, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19590.

Fraser:2018:RRL

[Fra18] Jennifer Fraser. Review: Rebecca Lemov. Database of Dreams: The Lost Quest to Catalog Humanity. 354pp. New Haven: Yale University Press, 2015. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):183–185, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27204.

Freiman:2016:RCR

[Fre16] Ori Freiman. Review: Cass R. Sunstein. Conspiracy Theories and Other Dangerous Ideas. 304 pp. Simon & Schuster, 2014. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):100–104, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/21013.

Fuller:2007:SDE

[Ful07] Steve Fuller. Science democratised = expertise decommissioned. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1): 25, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2969.

Fuller:2008:SSG

[Ful08] Steve Fuller. Science studies goes public: A report on an ongoing performance. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2 (1):11, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5069.

Fuller:2009:RL

[Ful09]  Steve Fuller.  Response to

Lynch. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):220– 222, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/8473.

Fuller:2011:RMT

[Ful11] Steve Fuller. A response to Mike Thicke (2011). Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):75–78, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/15341.

Fuller:2013:CPK

[Ful13] Steve Fuller. On commodification and the progress of knowledge in society: A defence. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):12–20, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/20075.

Gavrus:2010:RAC

[Gav10] Delia Gavrus. Review: Aaron A. Cohen-Gadol and Dennis D. Spencer. The Legacy of Harvey Cushing: Profiles of Patient Care. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):280– 282, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/13946.

Gelfert:2022:BPP

[Gel22] Axel Gelfert. Between pedantry and populism: a Kantian perspective on the assault on scientific expertise. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1): 113–122, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38207.

Gross:2012:VRS

[GL12] Ari Gross and Eleanor Louson. Visual representation and science: Editors’ introduction. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):1–7, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/18644.

Goldberg:2016:WTT

[Gol16] Daniel Goldberg. “What they think of the causes of so much suffering”: S. Weir Mitchell, John Kearsley Mitchell, and ideas about phantom limb pain in Late 19th c. America. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1): 27–54, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/20844.

Gooday:2009:DMS

[Goo09] Graeme Gooday. Domesticating the magnet: Secularity, secrecy and ‘permanency’ as epistemic boundaries in Marie Curie’s early work. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1): 68–81, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/10615.

Goodwin:2012:VRS

[Goo12] William Goodwin. Visual representations of structure and the dynamics of scientific modeling. Spontaneous Generations:

A Journal for the History and Philosophy of Science, 6(1):131– 141, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/17155.

Greene:2022:EMD

[Gre22] Mott Greene. Experts, managerialism, and democratic theory. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):1– 21, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38194.

Gil-Riano:2009:EIE

[GRH09] Sebastian Gil-Riano and Vivien Hamilton. Editor’s introduction: Epistemic boundaries. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):1–8, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/11182.

Gross:2009:RPT

[Gro09] Ari Gross. Review: Phillip Thurtle. The Emergence of Genetic Rationality: Space, Time, & Information in American Biological Science, 1870–1920. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):229–232,

????  2009.  CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5980.

Hacking:2009:RPB

[Hac09] Ian Hacking. Response to Professor Blute. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1): 226–228, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/10061.

Hallet:2009:SGH

[Hal09] Dani Hallet. On the subject of Goethe: Hermann von Helmholtz on Goethe and scientific objectivity. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):178– 194, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6568.

Hamilton:2009:RMR

[Ham09] Vivien Hamilton. Review: Maria Rentetzi. Trafficking Materials and Gendered Research Practices: Radium Research in Early 20th Century Vienna. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):246–248, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11315.

Harre:2010:EE

[Har10] Rom Harré. Equipment for an experiment. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 30–38, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/11334.

Hehmeyer:2010:CAS

[Heh10] Ingrid Hehmeyer. The challenge of authenticating scientific objects in museum collections: Exposing the forgery of a Moroccan astrolabe allegedly dated 1845 CE. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):8– 20, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/13025.

Hentschel:2012:SDS

[Hen12] Klaus Hentschel. The Stuttgart Database of Scientific Illustrators 1450–1950: Making the invisible hands visible. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 182–191, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/17156.

Hendry:2018:SRH

[Hen18] Robin Hendry. Scientific realism and the history of chemistry. Spontaneous Generations:

A Journal for the History and Philosophy of Science, 9(1):108– 117, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/28062.

Hooker:2018:MLF

[HH18] Giles Hooker and Cliff Hooker.

Machine learning and the future of realism. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 174–182, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27047.

Hooker:2022:EFO

[HHH22] Cliff Hooker, Claire Hooker, and Giles Hooker. Expertise, a framework for our most characteristic asset and most basic inequality. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):27– 35, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38196.

Hull:2007:EPC

[Hul07] James Hull. The expert professor: C. R. Young and the Toronto Building Code. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1 (1):86, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.

                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/2974.

Hull:2010:LFRa

[Hul10a] James Hull. Let freeness ring: The Canadian Standard Freeness Tester as hegemonic engine. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):61– 70, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/15962.

Hull:2010:LFRb

[Hul10b] James Hull. Let freeness ring: The Canadian Standard Freeness Tester as hegemonic engine. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):61– 70, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/11331.

Humphreys:2010:CSC

[Hum10] Paul Humphreys. Conceptual sea changes. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1): 111–115, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14185.

Jackson:2008:GMS

[Jac08] J. Kasi Jackson. Gender, mad scientists and nanotechnology. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):45, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/3516.

James:2008:ECC

[Jam08] Christine A. James. Evolution and conservative Christianity: How philosophy of science pedagogy can begin the conversation. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1): 185, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/3553.

Jones-Imhotep:2012:SV

[JI12] Edward Jones-Imhotep. Sound and vision. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 191–202, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16216.

John:2022:TVS

[Joh22a] Stephen John. The two virtues of science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1): 47–53, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38199.

Johnson:2022:RHH

[Joh22b] Bethany Johnson. Review of Hearing Happiness: Deafness Cures in History. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1): 125–128, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38210.

Kearnes:2008:IMS

[Kea08] Matthew Kearnes. Informationalising matter: Systems understandings of the nanoscale. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):99, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/5073.

Keller:2011:WCS

[Kel11] Evelyn Fox Keller. What are climate scientists to do? Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):19–26, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/15165.

Kemp:2012:TMO

[Kem12] Martin Kemp. “The testimony of my own eyes”: The strange case of the mammal with a beak. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):43– 49, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/17157.

Kulic:2010:RMS

[Kul10] Anthony Kulic. Review: Michael Strevens. Depth: An Account of Scientific Explanation. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):292–299, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14195.

Kusukawa:2012:TKC

[Kus12] Sachiko Kusukawa. Thomas Kirke’s copy of Philosophical Transactions. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 8–14, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16374.

Ladyman:2018:SRA

[Lad18] James Ladyman. Scientific realism again. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 99–107, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/29356.

Lander:2011:BTC

[Lan11] Bryn Lander. Between theory and craft: Exploring the role of co-operation within scientific research labs. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1): 58–74, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14957.

Lehoux:2009:CME

[Leh09] Daryn Lehoux. Creation myths and epistemic boundaries. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):28–34, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/8472.

Leifer:2016:RDS

[Lei16] Matthew Saul Leifer. Review: A. Douglas Stone. Einstein and the Quantum: The Quest of the Valiant Swabian. 332 pp. Princeton University Press, 2013. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):105–108, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/20940.

Levina:2009:EEB

[Lev09] Marina Levina. Exploring epistemic boundaries between scientific and popular cultures. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):105–112, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6569.

Levere:2010:AER

[Lev10] Trevor H. Levere. Apparatus and experimentation revisited. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):148–154, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11892.

Lewis:2012:RFG

[Lew12] Cory Lewis. Review: Frederick Grinnell, The Everyday Practice of Science: Where Intuition and Passion Meet Objectivity and Logic. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):242– 244, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16152.

Lightman:2011:PC

[Lig11]  Bernard Lightman.  Periodi-

cals and controversy. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):5–11,

????  2011.  CODEN ????

ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/15324.

Lorenz:2022:RDT

[Lor22]  Lissette Lorenz.  Review of

Deep Time Reckoning: How Future Thinking Can Help Earth Now. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):135– 137, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38214.

Lowrie:2012:AOH

[Low12] Ian Lowrie. On adaptive optics: The historical constitution of architectures for expert perception in astronomy. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 203–224, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/16133.

Lynch:2007:ESC

[Lyn07] Michael Lynch. Expertise, skepticism and cynicism: Lessons from science & technology studies. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):17, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/2968.

Lynch:2009:GPC

[Lyn09] Michael Lynch. Going public: A cautionary tale. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):213–219, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6085.

Lyons:2018:FCE

[Lyo18] Timothy D. Lyons. Four challenges to epistemic scientific realism — and the Socratic alternative. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):146– 150, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/26993.

Maienschein:2011:WDS

[Mai11] Jane Maienschein. Why do stem cells create such public controversy? Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):27– 35, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/15099.

McCullough:2010:OCR

[McC10] Philip Murray McCullough. Otto in the Chinese Room. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):129–137,

????  2010.  CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11718.

McKenzie:2018:BRC

[McK18] Kerry McKenzie. Being realistic: the challenge of theory change for a metaphysics of scientific realism. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):136– 142, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/26998.

Menon:2016:RRD

[Men16] Alka Vaid Menon. Review: Roberts, Dorothy. 2011. Fatal Invention: How Science, Politics, and Big Business Re-Create Race in the Twenty-First Century. New York: The New Press. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):109–111, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/20388.

Mahony:2012:CRE

[MH12] Martin Mahony and Mike Hulme. The colour of risk: An exploration of the IPCC’s “burning embers” diagram. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 75–89, ???? 2012. CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16075.

Miller:2007:SEE

[Mil07a] Boaz Miller. Scientific expertise:

Epistemological worries, political dilemmas (Focused Discussion Editor’s introduction). Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):13, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2967.

Miller:2007:WTS

[Mil07b] Boaz Miller. What trust in science? Review of the trust in science workshop. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1 (1):132, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2979.

Milburn:2008:AAV

[Mil08]  Colin Milburn.  Atoms and

avatars: Virtual worlds as massively-multiplayer laboratories. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):63, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/4895.

Miller:2011:RLM

[Mil11]  Boaz Miller.  Review: Lee

McIntyre. Dark Ages: The Case for a Science of Human Behavior. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):85– 87, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14576.

Martin-Nielsen:2007:EVI

[MN07] Janet Martin-Nielsen. An engineer’s view of an ideal society: The economic reforms of C. H. Douglas, 1916–1920. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):95, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/978.

Moore:2010:RAP

[Moo10] R. Moore. Review: Adrian Parr. Hijacking Sustainability. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):283–285, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14196.

Morris:2009:RCF

[Mor09] Kevin Morris. Review: Carl F. Craver. Explaining the Brain: Mechanisms and the Mosaic Unity of Neuroscience. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):239–242, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6126.

Musgrave:2018:BMD

[Mus18] Alan Musgrave. BEWARE of Mad DOG Realist. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 52–64, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27051.

Martin:2007:MPE

[MW07] Aaron K. Martin and Edgar A.

Whitley. Managing public expectations of technological systems: A case study of a problematic government project. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):67, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2973.

Myburgh:2022:STU

[Myb22] Brittany Myburgh. Space-time and utopia notes on artistic engagement with physics from cubism to Laszlo Moholy-Nagy. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):54–62,

????  2022.  CODEN ????

ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38200.

Myers:2008:CML

[Mye08] Natasha Myers. Conjuring machinic life. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1): 112, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/4894.

Naylor:2022:HBE

[Nay22] Robert Naylor. How (not) to build an expert: Personal reflections on a terminated physics career. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):98– 106, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38203.

Niemoczynski:2009:PAV

[Nie09] Leon Niemoczynski. Phenomenology in the American vein: Justus Buchler’s ordinal naturalism and its importance for the justification of epistemic objects. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):9–27, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/6577.

Nugayev:2016:MEG

[Nug16] Rinat Magdievich Nugayev. Maxwellian electrodynamics genesis and development: Intertheoretic context. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):55– 92, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19467.

Oakes:2018:RDV

[Oak18a] Andrew Oakes. Review: Douglas A. Vakoch and Matthew F. Dowd. The Drake Equation: Estimating the Prevalence of Extraterrestrial Life through the Ages. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):186– 188, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/29357.

Oakes:2018:RMB

[Oak18b] Andrew Oakes. Review: Melinda Baldwin. Making Nature: The History of a Scientific Journal. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):189– 191, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/29358.

Obrist:2012:VRS

[Obr12] Barbara Obrist. Visual representation and science: Visual figures of the universe between antiquity and the Early Thirteenth Century. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 15–23, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/17158.

Olsen:2010:OTN

[Ols10] Dorothy Sutherland Olsen. “Old” technology in new hands: Instruments as mediators of interdisciplinary learning in microfluidics. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):231– 254, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11942.

Ottinger:2009:EFA

[Ott09] Gwen Ottinger. Epistemic fencelines: Air monitoring instruments and expert-resident boundaries. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):55– 67, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6115.

Owens:2016:CSP

[Owe16] Kellie Owens. Colorblind science?: Perceptions of the importance of racial diversity in science research. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1): 13–21, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/20893.

Parker:2010:IWD

[Par10] Wendy S. Parker. An instrument for what? Digital computers, simulation and scientific practice. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):39– 44, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/13765.

Patterson:2010:RDP

[Pat10] Sarah-Jane Patterson. Review:

David Pantalony. Altered Sensations: Rudolph Koenig’s Acoustical Workshop in Nineteenth-Century Paris. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):289–291, ???? 2010. CO-DEN ???? ISSN 1913-0465. URL

https://spontaneousgenerations.
library.utoronto.ca/index.php/
SpontaneousGenerations/article/
view/14194.

Perini:2012:TBT

[Per12]  Laura Perini.  Truth-bearers

or truth-makers? Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):142–147,

????  2012.  CODEN ????

ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/17357.

Petrie:2010:RWS

[Pet10] Bruce J. Petrie. Review: William Sims Bainbridge. The Warcraft Civilization: Social Science in a Virtual World. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):270–272, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/13757.

Peterson:2011:RSP

[Pet11]  Andrew Peterson.  The rele-

vance of scientific practice to the problem of coordination. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):44–57, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14607.

Pitt:2008:STN

[Pit08] Joseph C. Pitt. Small talk: Nanotechnology and metaphor. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):90, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/3341.

Pooley:2022:RIM

[Poo22] Jeff Pooley. Review of The Information Manifold: Why Computers Can’t Solve Algorithmic Bias and Fake News. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1): 138–139, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38218.

Price:2007:STC

[Pri07] Huw Price. Starving the theological cuckoo: Review of John Leslie. Infinite Minds: A Philosophical Cosmology. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1): 136, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/2980.

Psillos:2018:TAR

[Psi18] Stathis Psillos. Tolstoy’s argument: realism and the history of science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 68–77, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/28059.

Pursell:2016:TSI

[Pur16] Caroll Pursell. Technology and social inequality. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):22–26, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/22610.

Ranalli:2011:PPR

[Ran11] Brent Tibor Ranalli. A prehistory of peer review: Religious blueprints from the Hartlib Circle. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):12– 18, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14973.

Record:2008:FLS

[Rec08] Isaac Record. Frankenstein in Lilliput: Science at the nanoscale (Editor’s introduction). Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):22, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5070.

Record:2009:RDR

[Rec09] Isaac Record. Review: Daniel Rothbart. Philosophical Instruments: Minds and Tools at Work. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):233– 235, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.

utoronto.ca/index.php/SpontaneousGenerations/
article/view/6086.

Record:2010:SIK

[Rec10] Isaac Record. Scientific instruments: Knowledge, practice, and culture [Editor’s introduction]. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):1– 7, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14231.

Rehman:2022:PTJ

[Reh22] Rashad Rehman. Philosophizing together: Josef Pieper on philosophical responsibility and expertise. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):88– 97, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38202.

Rietmann:2011:WHR

[Rie11] Felix Ernst Rietmann. “What is HPS for?” Review of the Fifth Joint Workshop on Integrated History and Philosophy of Science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1): 88–90, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14189.

Record:2008:RPC

[RM08] Isaac Record and Andrew Munro. Review: Paul E. Ceruzzi. Internet Alley: High Technology in Tysons Corner, 1945–2005. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):251, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/5068.

Robinson:2016:RPS

[Rob16] Joan H. Robinson. Review of peoples’ science: Bodies and rights on the stem cell frontier. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):112–114, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/20538.

Ross:2007:WCA

[Ros07] Sage Rogers Ross. We cannot allow a Wikipedia gap! Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):1, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/1017.

Ross:2022:HTP

[Ros22] Christian Ross. Handservant of technocracy public engagement and expertise in heritable human genome editing. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):63–87, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38201.

Rouse:2018:BRA

[Rou18] Joseph Rouse. Beyond realism and antirealism — at last? Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):46–51, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/26979.

Rozell:2013:RPF

[Roz13] Daniel Rozell. Review: Paul Feyerabend, The Tyranny of Science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):97– 99, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19247.

Rowland:2011:RBL

[RPK11] Nicholas J. Rowland, Jan-Hendrik Passoth, and Alexander B. Kinney. Review: Bruno Latour. Reassembling the Social: An introduction to Actor-Network-Theory. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):95– 99, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.

                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/14968.

Sankey:2018:DSR

[San18] Howard Sankey. A dilemma for the scientific realist. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):65–67, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/26352.

Scharf:2007:RGC

[Sch07] Sara Scharf. Review: Geoffrey C. Bowker. Memory Practices in the Sciences. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1 (1):149, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/2982.

Scharf:2008:MII

[Sch08a] Sara Scharf. Multiple independent inventions of a non-functional technology: Combinatorial descriptive names in botany, 1640– 1830. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):145, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/3552.

Scharf:2008:RHC

[Sch08b] Sara Scharf. Review: Hine, Christine. 2008. Systematics as Cyberscience. Cambridge, MA: MIT Press. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1): 245, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/4852.

Schummer:2008:PET

[Sch08c] Joachim Schummer. The popularization of emerging technologies through ethics: From nanotechnology to synthetic biology. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):56, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/3529.

Schranz:2018:RAH

[Sch18] Kristen M. Schranz. Review: Andre Holenstein, Hubert Steinke, and Martin Stuber, eds. Scholars in Action: The Practice of Knowledge and the Figure of the Savant in the 18th Century, Volumes 1 and 2. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1): 192–195, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/27555.

Schrempp:2022:BPS

[Sch22] Gregory Schrempp. The best popular science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1): 22–26, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38179.

Sent:2013:ESH

[Sen13] Esther-Mirjam Sent. The economics of science in historical and disciplinary perspective. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):6–11, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/19624.

Shaw:2018:WRD

[Sha18] Jamie Craig Owen Shaw. Why the realism debate matters for science policy: The case of the Human Brain Project. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):82–98, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/27760.

Shindell:2010:DPI

[Shi10] Matthew Benjamin Shindell. Domesticating the planets: Instruments and practices in the development of planetary geology. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):191– 230, ???? 2010. CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11954.

Sismondo:2022:RST

[Sis22]  Sergio Sismondo.  Review of

Strange Trips: Science, Culture, and the Regulation of Drugs. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):123– 124, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38209.

Sewell:2010:MCD

[SJ10] Jessica Ellen Sewell and Andrew Johnston. Material culture and the Dobsonian telescope. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):155– 162, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11947.

Suay-Matallana:2012:VRS

[SMCL12] Ignacio Suay-Matallana and Mar Cuenca-Lorente. “Visual representations in science”: Review of the 6th European Spring School on History of Science and Popularization: International Workshop, May 19–21 2011, Maó, Menorca, Spain. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):245– 251, ???? 2012. CODEN ????

ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/16159.

Spencer:2012:TIC

[Spe12] Matt Spencer. Trouble with images in computational physics. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):34–42, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/16119.

Stanford:2018:FFA

[Sta18] P. Kyle Stanford. A fond farewell to ‘approximate truth’? Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):78–81, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/28057.

Stelmackowich:2012:ICD

[Ste12] Cindy Stelmackowich. The instructive corpse: Dissection, anatomical specimens, and illustration in Early Nineteenth-Century medical education. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):50–64, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/17159.

Stephens:2022:RWT

[Ste22] Christopher Stephens. Review of Why Trust Science? Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):140– 144, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/38220.

Stuart:2012:RJR

[Stu12]  Michael T. Stuart.  Review:

James R. Brown, Laboratory of the Mind. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1): 237–241, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/15820.

Taylor:2012:H

[Tay12] Bruce Taylor. Holdings. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):233–236, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/17362.

Turner:2013:BEC

[TDFM13] Andrew Turner, Clara Dallaire-Fortier, and Madeleine J. Murtagh. Biobank economics and the “commercialization problem”. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):69–80,

????  2013.  CODEN ????

ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19555.

Teller:2018:RPR

[Tel18]  Paul Teller.  Referential and

perspectival realism. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):151–164, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/26990.

Theodore:2009:WKM

[The09] David Theodore. Was Kekulé’s mind brainbound? The historiography of chemistry and the philosophy of extended cognition. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):158– 177, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6125.

Thicke:2011:RSF

[Thi11] Mike Thicke. Review: Steve Fuller. Science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5 (1):91–94, ???? 2011. CODEN ???? ISSN 1913-0465. URL

https://spontaneousgenerations.
library.utoronto.ca/index.php/
SpontaneousGenerations/article/
view/14294.

Thicke:2013:EAS

[Thi13a] Mike Thicke. Economic aspects of science: Editor’s introduction. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):1– 5, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/20379.

Thicke:2013:RDT

[Thi13b] Mike Thicke. Review: David Tyfield, The Economics of Science: A Critical Realist Overview, Volumes 1 and 2. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1): 94–96, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/18795.

Thicke:2016:RCW

[Thi16] Mike Thicke. Review: Cold War Social Science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1):115–117, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19825.

Thoreau:2010:TPM

[TN10] François Thoreau and Maria Neicu. Taming the “publication machine”: Generating unity, engaging the trading zones. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):163–172, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/11944.

Tedre:2009:CNM

[TS09] Matti Tedre and Erkki Sutinen. Crossing the Newton–Maxwell gap: Convergences and contingencies. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):195– 212, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/3388.

Tucker:2012:HWS

[Tuc12] Jennifer Tucker. “The hidden world of science”: Nature as art in 1930’s American print advertising. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):90– 105, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/17160.

Turner:2007:PEE

[Tur07] Stephen Turner. Political epistemology, experts, and the aggregation of knowledge. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1):36, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.

utoronto.ca/index.php/SpontaneousGenerations/
article/view/2970.

Turrini:2013:CVM

[Tur13] Mauro Turrini. Computer versus microscope: Visual activity fields of instruments in the information age. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1): 81–93, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16151.

Tyfield:2013:TSR

[Tyf13] David Tyfield. Transition to Science 2.0: “Remoralizing” the economy of science. Spontaneous Generations: A Journal for the History and Philosophy of Science, 7(1):29–48, ???? 2013. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/19664.

Varughese:2011:MSD

[Var11] Shiju Sam Varughese. Media and science in disaster contexts: Deliberations on earthquakes in the regional press in Kerala, India. Spontaneous Generations: A Journal for the History and Philosophy of Science, 5(1):36– 43, ???? 2011. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/14969.

Vickers:2018:QVS

[Vic18] Peter Vickers. Quo vadis selective scientific realism? Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):118–121, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/28056.

Virdi:2008:BKG

[Vir08a] Jaipreet Virdi. Bridging the knowledge gap: Examining potential limits in nanomedicine. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):25, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/3517.

Virdi:2008:RFA

[Vir08b] Jaipreet Virdi. Review: Fritz Allhoff and Patrick Lin, Eds. Nanotechnology and Society: Current and Emerging Ethical Issues. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):248, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/5071.

Virdi:2010:LAR

[Vir10] Jaipreet Virdi. Learning from artifacts: A review of the “Reading Artifacts: Summer Institute in the Material Culture of Science,” presented by The Canada Science and Technology Museum and Situating Science Cluster. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):276–279, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/13861.

Viseu:2008:DN

[Vis08] Ana Viseu. Disciplining nano.

Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):122, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/5074.

vanTol:2007:MBC

[vT07] Deanne van Tol. Mothers, babies, and the colonial state: The introduction of maternal and infant welfare services in Nigeria, 1925–1945. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1): 110, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/1761.

vanWeelden:2022:REI

[vW22] Joseph van Weelden. Review of Expertise: an Introduction. Spontaneous Generations: A Journal for the History and Philosophy of Science, 10(1):145– 148, ???? 2022. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/38221.

Wallace:2007:AOE

[Wal07] Stephen Wallace. Ann Oakley. Experiments in knowing: Gender and method in the social sciences. Spontaneous Generations: A Journal for the History and Philosophy of Science, 1(1): 151, ???? 2007. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/2983.

Walsh:2008:CBU

[Wal08] Denis Walsh. A commentary on Blute’s ‘updated definition’. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1): 6, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/4931.

Wald:2009:DDD

[Wal09] Benjamin Elliott Wald. Dealing with disagreement: Distinguishing two types of epistemic peers. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):113– 122, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

                                        spontaneousgenerations.library.
                                        utoronto.ca/index.php/SpontaneousGenerations/
                                        article/view/6562.

Weidenhammer:2010:DSH

[WD10] Erich Weidenhammer and Michael Da Silva. Out the door: A short history of the University of Toronto Collection of Historical Scientific Instruments. Spontaneous Generations: A Journal for the History and Philosophy of Science, 4(1):255–261, ???? 2010. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/13802.

Westmoreland:2009:UBE

[Wes09] Mark William Westmoreland. The unbounded bridge of emergent evolution, a review of Jacob Klapwijk, Purpose in the Living World: Creation and Emergent Evolution. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):243– 245, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6105.

Winsberg:2008:MTN

[Win08] Eric Winsberg. Models and theories at the nano-scale. Spontaneous Generations: A Journal for the History and Philosophy of Science, 2(1):139, ???? 2008. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/4840.

Wray:2018:REH

[Wra18] K. Brad Wray. The relevance of evidence from the history of science in the contemporary realism/anti-realism debate. Spontaneous Generations: A Journal for the History and Philosophy of Science, 9(1):143–145, ???? 2018. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/26986.

Wright:2009:HEP

[Wri09] Aaron Sidney Wright. “I hold every properly qualified navigator to be a philosopher”: The making of the U.S. Naval Observatory’s Global Laboratory. Spontaneous Generations: A Journal for the History and Philosophy of Science, 3(1):82–94, ???? 2009. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/6116.

Wuthrich:2012:IFD

[Wüt12] Adrian Wüthrich. Interpreting Feynman diagrams as visual models. Spontaneous Generations: A Journal for the History and Philosophy of Science, 6(1):172– 181, ???? 2012. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/16131.

Wood:2016:SSI

[WW16] Christine V. Wood and Simon N. Williams. STS and social inequality: Editor’s introduction. Spontaneous Generations: A Journal for the History and Philosophy of Science, 8(1): 1–2, ???? 2016. CODEN ???? ISSN 1913-0465. URL https://

spontaneousgenerations.library.
utoronto.ca/index.php/SpontaneousGenerations/
article/view/26686.